### v2

# A Note on the Phragmén-Lindelöf Theorem

Abstract.
We provide a generalization of the Phragmén-Lindelöf principal of Rademacher with the aim of correcting, or at least provide a pathway to correcting, several errors appearing in the literature.
Andrew Fiori thanks and acknowledges the University of Lethbridge for their financial support as well as the support of NSERC Discovery Grant RGPIN-2020-05316.
The Phragmén-Lindelöf principle is well known and well used in analytic number theory as a tool to obtain bounds on $L$-functions by interpolating bounds obtainable on vertical lines. One of the early general purpose results which is now widely used is [Rademacher, Theorem 2]. There is a generalization of this result presented in [Tr14-2, Lemma 3]. The proof of Rademacher is quite technical as it requires the construction of an analytic function with very precisely described growth properties. The proof of Trudgian has an error111the function $F(s)$ defined in the proof is not in general analytic as it is a function of $\sigma={\rm Re}(s)$. which arises in trying to avoid doing the same technical argument with power of $\log(Q+s)$ rather than $(Q+s)$.
In what follows we sufficiently correct the error so that proofs using it should be recoverable, at worst at the expense of some constants. Moreover, we provide a generalized version Rademacher’s work using what is in our view a simpler proof than that of Rademacher. We note that the result of Trudgian has been widely cited (45 direct citations, many of which have even more), including several results which now improve upon Trudgian’s main result (see [HSW, BW]). Consequently, that our result in practice recovers (at the expense of extra checks for small values, see Section 1) and improves upon Trudgian’s work (see Sections 2 and 3) means our work shall either recover these other results, or at least sets the groundwork for others to do so.
Theorem 1 (Phragmén-Lindelöf principle).
Let $a,b$ be real numbers, and let $G_{1},\ldots,G_{r}$ be complex functions that are holomorphic for ${\rm Re}(s)\in[a,b]$, have $\overline{G(s)}=G(\overline{s})$, and satisfy the growth condition
$C_{1}\exp\left(-C_{2}e^{C_{3}|t|}\right)<|G_{i}(\sigma+it)|$
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$ for $\sigma\in[a,b]$. Suppose that $f$ is a holomorphic function on $a\leq\sigma\leq b$ and satisfies the growth condition
$|f(\sigma+it)|<C_{1}\exp\left(C_{2}e^{C_{3}|t|}\right)$
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$. Suppose that we have real numbers $\alpha_{1},\ldots,\alpha_{r}\geq 0$ and $\beta_{1},\ldots,\beta_{r}\geq 0$ such that
$|f(a+it)|\leq\bigg{|}\prod_{i=1}^{r}G_{i}(a+it)^{\alpha_{i}}\bigg{|}$
and
$|f(a+it)|\leq\bigg{|}\prod_{i=1}^{r}G_{i}(b+it)^{\beta_{i}}\bigg{|}.$
Finally, suppose $|G_{i}(\sigma+it)|$ is increasing in $\sigma$ if $\alpha_{i}>\beta_{i}$ and decreasing if $\beta_{i}<\alpha_{i}$. Then for $a\leq\sigma\leq b$, we have
$|f(\sigma+it)|\leq\bigg{|}\prod_{i=1}^{r}G_{i}(\sigma+it)^{\alpha_{i}}\bigg{|}^{\frac{b-\sigma}{b-a}}\bigg{|}\prod_{i=1}^{r}G_{i}(\sigma+it)^{\beta_{i}}\bigg{|}^{\frac{\sigma-a}{b-a}}.$
Remark 2.
The result can be extended to have $\alpha_{i}$ and $\beta_{i}$ negative at the expense of a growth condition
$|G_{i}(\sigma+it)|<C_{1}\exp\left(C_{2}e^{{C_{3}}|t|}\right)$
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$ for $\sigma\in[a,b]$ by simply multiplying $f(s)$ by $G_{i}(s)^{-\min(\alpha_{i},\beta_{i})}$.
Some standard choices for $G_{i}(s)$ are
  1. (1)
    $G(s)=e$ where $e^{\alpha}=A$ and $e^{\beta}=B$.
  2. (2)
    $G(s)=(Q+s)$ where $Q>-a$.
Corollary 3.
Let $a$ and $b$ be real numbers and suppose $f(s)$ is holomorphic for ${\rm Re}(s)\in[a,b]$ and satisfies the growth condition
$|f(\sigma+it)|<C_{1}\exp\left(C_{2}e^{{C_{3}}|t|}\right)$
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$.
Suppose there exist $A,B,\alpha_{1}\geq\beta_{1},\alpha_{2}=\beta_{2},\alpha_{3}=\beta_{3},Q_{1}>-a,Q_{2}>1-a,Q_{3}>e-a$ such that:
$|f(a+it)|\leq A\big{|}Q_{1}+a+it\big{|}^{\alpha_{1}}\big{|}\log(Q_{2}+a+it)\big{|}^{\alpha_{2}}\big{|}\log\log(Q_{3}+a+it)\big{|}^{\alpha_{3}}$
and
$|f(b+it)|\leq B\big{|}Q_{1}+b+it\big{|}^{\beta_{1}}\big{|}\log(Q_{2}+b+it)\big{|}^{\beta_{2}}\big{|}\log\log(Q_{3}+b+it)\big{|}^{\beta_{3}}$
then if $\sigma\in[a,b]$ we have $|f(\sigma+it)|$ is bounded by
$\bigg{(}A\big{|}Q_{1}$ $+\sigma+it\big{|}^{\alpha_{1}}\big{|}\log(Q_{2}+\sigma+it)\big{|}^{\alpha_{2}}\big{|}\log\log(Q_{3}+\sigma+it)\big{|}^{\alpha_{3}}\bigg{)}^{\frac{b-\sigma}{b-a}}\cdot$
$\bigg{(}B\big{|}Q_{1}+\sigma+it\big{|}^{\beta_{1}}\big{|}\log(Q_{2}+\sigma+it)\big{|}^{\beta_{2}}\big{|}\log\log(Q_{3}+\sigma+it)\big{|}^{\beta_{3}}\bigg{)}^{\frac{\sigma-a}{b-a}}.$
Proof.
The functions $G_{1}(s)=e$, $G_{2}(s)=Q_{1}+s$, $G_{3}(s)=\log(Q_{2}+s)$ and $G_{4}(s)=\log\log(Q_{3}+s)$ will satisfy all the hypotheses of the theorem given that $\alpha_{2}=\beta_{2}$ and $\alpha_{3}=\beta_{3}$. ∎
Remark 4.
The function $|\log Q+s|$ is neither always increasing or always decreasing in $\sigma$. The behavior depends on $t$ and $\sigma$. For this reason the argument we gave above only works with $\alpha_{2}=\beta_{2}$ and $\alpha_{3}=\beta_{3}$.
Moreover, you should note well that $|\log(Q_{2}+s)|\geq\log|Q_{2}+s|$ and $|\log\log(Q_{2}+s)|\geq\log\log|Q_{2}+s|$. Indeed
$|\log(Q_{2}+s)|=\bigg{(}1+\frac{{\rm arg}(Q_{2}+s)^{2}}{(\log|Q_{2}+s|)^{2}}\bigg{)}^{1/2}(\log|Q_{2}+s|)$
and
$|\log\log(Q_{2}+s)|=\bigg{(}1+\frac{{\rm arg}(\log(Q_{2}+s))^{2})}{(\log|\log(Q_{2}+s)|)^{2}}\bigg{)}^{1/2}\log|\log(Q_{2}+s)|$
where
${\rm arg}(\log(Q_{2}+s))=\arctan\bigg{(}\frac{{\rm arg}(Q_{2}+s)}{\log|Q_{2}+s|}\bigg{)}.$
This difference can be relevant in both the hypothesis and conclusion of the theorem depending on the signs of the exponents. There are multiple results in the literature which fail to account for this. See Sections 2 and 3 for a method to improve this.
Lemma 5.
With the hypothesis as in Theorem 1 we have
$|f((a+b)/2+it)|<\left|\prod_{i=1}^{r}G_{i}((a+b)/2+it)^{\alpha_{i}+\beta_{i}}\right|^{1/2}.$
That is, Theorem 1 holds where $\sigma$ is the midpoint between $a$ and $b$.
Proof.
We shall denote by $\overline{f}$ the analytic function $\overline{f}(s)=\overline{f(\overline{s})}$. We note that a bound on $|f(s)|$ in terms of $G_{i}(s)$ satisfying $\overline{G_{i}(s)}=G_{i}(\overline{s})$ implies the same bound on $|\overline{f}(s)|$.
Consider the function
$F(s)=f(a+s)\overline{f}(b-s).$
Define
$H_{i}(s)=\begin{cases}G_{i}(a+s)^{\alpha_{i}}G_{i}(b-s)^{\beta_{i}}&\alpha_{i}>\beta_{i}\text{ and }G(a+\sigma+it)\text{ increasing for all }t\\ G_{i}(a+s)^{\beta_{i}}G_{i}(b-s)^{\alpha_{i}}&\alpha_{i}<\beta_{i}\text{ and }G(a+\sigma+it)\text{ decreasing for all }t\\ G_{i}(a+s)^{\alpha_{i}}G_{i}(b-s)^{\beta_{i}}&\alpha_{i}=\beta_{i}\end{cases}$
and
$H(s)=\prod_{i=1}^{r}H_{i}(s)$
By our assumptions we have
$F(s)/H(s)$
is analytic in $[0,(b-a)]$ and it satisfies the necessary growth condition to apply the standard Phragmén-Lindelöf principle (see for example [Rudin, Theorem 12.9]). Moreover, for ${\rm Re}(s)=0$ and ${\rm Re}(s)=(b-a)$ we have have $F(s)/H(s)\leq 1$ as
$F(it)=|f(a+it)\overline{f}(b-it)|=|f(a+it)f(b+it)|$
and
$|F(((b-a)+it)|=|f(b+it)\overline{f}(a-it)|=|f(a+it)f(b+it)|$
are from our hypotheses bounded by
$\prod_{i=1}^{r}|G_{i}(a+it)|^{\alpha_{i}}|G_{i}(b+it)|^{\beta_{i}}.$
Now notice that $|H_{i}(it)|$ and $|H_{i}((b-a)+it)|$ take on the values222which is taken on at which point depends on the case we are considering.
$|G_{i}(a+it)|^{\alpha_{i}}|G_{i}(b+it)|^{\beta_{i}}\qquad\text{and}\qquad|G_{i}(a+it)|^{\beta_{i}}|G_{i}(b+it)|^{\alpha_{i}}.$
Under the hypotheses the second is larger than the first and the first is the needed bound.
Thus by the Phragmén-Lindelöf principle we have $F(s)/H(s)\leq 1$ on the whole interval. It follows by evaluating at $s=(b-a)/2+it$ that for all $t$ we have
$|F((a-b)/2+it)|\leq|H((a-b)/2+it)|=\prod_{i=1}^{r}|G_{i}((a+b)/2+it)|^{\alpha_{i}+\beta_{i}}.$
Since
$|F((a-b)/2+it)|=|f((a+b)/2+it)\overline{f}((a+b)/2-it)|=|f((a+b)/2+it)|^{2}$
we obtain the result. ∎
Proof of Theorem 1.
We first claim that the result holds for all $\sigma$ of the form
$a+(b-a)\frac{n}{2^{k}}$
where $\frac{n}{2^{k}}\in[0,1]$. The proof of this proceeds by induction on $k$, using Lemma 5 applied to the interval $a=\tfrac{\lfloor n/2\rfloor}{2^{k-1}}$ and $b=\tfrac{\lceil{n/2}\rceil}{2^{k-1}}$ with the bounds obtained from the inductive hypothesis.
The final result now follows by the continuity of $f(s)$ and $G_{i}(s)$. ∎
Remark 6.
We note that in most applications as functions of $t$ we will have $|G_{i}(a+it)|\sim|G_{i}(\sigma+it)|\sim|G_{i}(b+it)|$ so that if one wants to work with a function where $G_{i}$ which is decreasing (respective increasing) in $\sigma$ rather then increasing (respective decreasing) one can replace $G_{i}(s)$ by $G(b-a-s)$ at the expense of possibly needing to multiply by a constant to preserve the inequalities on the boundary.
1. Dealing with the Condition of Increasing/Decreasing in $\sigma$
Even though the function $|\log Q+s|$ is neither always increasing or always decreasing in $\sigma$ it is desirable to be able to use it as $G_{i}(s)$. There are options for dealing with such functions
  • Replace $G(s)$ with $G(Q^{\prime}-s)$ provided the function was at least consistently always increasing or decreasing. This is essentially what occurs in [Rademacher, Theorem 2] for situations $\beta>\alpha$ applied to $G(s)=Q+s$.
  • Replace $G(s)$ with $e^{c(s-a)}G(s)$ for some choice $c$ such that the resulting function is appropriately increasing/decreasing.
  • Grouping together several $G_{i}(s)$ so that the product satisfies the conditions.
  • Satisfy the increasing condition for $t>T_{0}$ and impose an alternative condition for $t<T_{0}$.
In each case we risk making the final result worse by a constant in order to satisfy the conditions. We provide a precise version of the last option below.
Theorem 7.
Let $a,b$ be real numbers, and let $G_{1},\ldots,G_{r}$ be complex functions that are holomorphic for ${\rm Re}(s)\in[a,b]$, have $\overline{G(s)}=G(\overline{s})$, and satisfy the growth condition
$C_{1}\exp\left(-C_{2}e^{{C_{3}}|t|}\right)<|G_{i}(\sigma+it)|$
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$ for $\sigma\in[a,b]$. Suppose that $f$ is a holomorphic function on $a\leq\sigma\leq b$ and satisfies the growth condition
$|f(\sigma+it)|<C_{1}\exp\left(C_{2}e^{{C_{3}}|t|}\right)$
for some $C_{1},C_{2}>0$ and $0<C_{3}<\frac{\pi}{b-a}$. Suppose that we have real numbers $\alpha_{1},\ldots,\alpha_{r}\geq 0$ and $\beta_{1},\ldots,\beta_{r}\geq 0$ such that
$|f(a+it)|\leq\bigg{|}\prod_{i=1}^{r}G_{i}(a+it)^{\alpha_{i}}\bigg{|}$
and
$|f(a+it)|\leq\bigg{|}\prod_{i=1}^{r}G_{i}(b+it)^{\beta_{i}}\bigg{|}.$
Suppose that for $|t|>T_{0}$ that $|G_{i}(\sigma+it)|$ is increasing in $\sigma$ if $\alpha_{i}>\beta_{i}$ and decreasing if $\beta_{i}<\alpha_{i}$. Finally, suppose that for each $t\leq T_{0}$ we have
$|f(\sigma+it)|<\prod_{i}{\rm min}_{\sigma}(|G(\sigma+it)|)^{\alpha_{i}\frac{b-\sigma}{b-a}+\beta_{i}\frac{\sigma-a}{b-a}}.$
Then for $a\leq\sigma\leq b$, we have
$|f(\sigma+it)|\leq\bigg{|}\prod_{i=1}^{r}G_{i}(\sigma+it)^{\alpha_{i}}\bigg{|}^{\frac{b-\sigma}{b-a}}\bigg{|}\prod_{i=1}^{r}G_{i}(\sigma+it)^{\beta_{i}}\bigg{|}^{\frac{\sigma-a}{b-a}}.$
Proof.
The proof is as before, except, to obtain the that
$F(s)/H(s)$
is bounded on the boundary for $|t|<T_{0}$ we must use the stronger hypothesis. We note that the hypothesis is strong enough to descend to the inductive cases which need to be considered. ∎
Remark 8.
To verify the condition
$|f(\sigma+it)|<\prod_{i}{\rm min}_{\sigma}(|G(\sigma+it)|)^{\alpha_{i}\frac{b-\sigma}{b-a}+\beta_{i}\frac{\sigma-a}{b-a}}$
one could check numerically on the whole region, alternatively, one can prove something stronger by using the maximum modulos condition to verify something on only the boundary.
Given that in practice one is interested in large $t$ one can typically arrange for the condition for small values to be satisfied by adjusting $G$ and preserving the asymptotics as one does in selecting $Q_{1}$ and $Q_{2}$ in Theorem 7.
2. Errors in the literature
There are unfortunately a number of errors in the literature where $|\log Q+s|$ was incorrectly handled as $\log|Q+s|$. So beyond the need to check the hypothesis which we needed in our corrected version the previous section would not allow one to fully recover the numerics of such results. In this section we highlight how one can potentially almost recover the numerics from such as error, at least for $t$ that are not too small.
Moreover, in number theory it is the case that bounds are often proven in the form
$|\zeta(1+it)|<C\log|t|\qquad t>e$
so that the hypothesis we have is actually stronger than a bound on $|\log\sigma+it|$. It would be desirable that the conclusions we prove would be similarly strong. As noted before the discrepancy here is that as $t\to\infty$
$|\log\sigma+it|=(1+O_{\sigma}(1/(\log|t|)^{2}))\log|t|.$
Unfortunately, the $(1/\log|t|)^{2}$ term shrinks slowly enough to impact numerics in some applications. Noticing that
$\tfrac{1}{2}\left|\log(1-(\sigma+it)^{2})\right|=(1+O_{\sigma}(1/|t|^{2})))\log|t|$
provides the core idea to improve this discrepancy. The essential idea here is that ${\rm arg}(1+\sigma+it)\sim-{\rm arg}(1-\sigma-it)$ so that we can obtain cancellation of error terms.
Now, given an interval $[a,b]$ we define functions
$G_{Q_{1},Q_{2}}(s)=\frac{1}{2}\log((Q_{1}+s)(Q_{2}-s))$
where $Q_{1}+a\geq 1$ and $Q_{2}-b\geq 1$. Notice that we have
$\log(Q_{1}+\sigma+it)$ $=\frac{1}{2}\log((Q_{1}+\sigma)^{2}+t^{2})+i\arctan\left(\frac{t}{Q_{1}+\sigma}\right)$
$=\log t+\log\left(1+\left(\frac{Q_{1}+\sigma}{t}\right)^{2}\right)+i\arctan\left(\frac{t}{Q_{1}+\sigma}\right)$
and
$\log(Q_{2}-\sigma-it)$ $=\frac{1}{2}\log((Q_{2}-\sigma)^{2}+t^{2})-i\arctan\left(\frac{t}{Q_{2}-\sigma}\right)$
$=\log t+\log\left(1+\left(\frac{Q_{2}-\sigma}{t}\right)^{2}\right)-i\arctan\left(\frac{t}{Q_{2}-\sigma}\right)$
so that with we have [^3] Using $\arctan(x/y)=\pi/2-\arctan{y/x}$.
$G_{Q_{1},Q_{2}}(\sigma+it)$ $=\log t+\log\left(1+\left(\frac{Q_{1}+\sigma}{t}\right)^{2}\right)\left(1+\left(\frac{Q_{2}-\sigma}{t}\right)^{2}\right)$
$\quad+i\left(\arctan\left(\frac{Q_{2}-\sigma}{t}\right)-\arctan\left(\frac{Q_{1}+\sigma}{t}\right)\right).$
Approximations for $\log$ and $\arctan$, which can be made precise, will give for $t>\!\!>0$ that
$\log\left(1+\left(\frac{Q_{1}+\sigma}{t}\right)^{2}\right)\left(1+\left(\frac{Q_{2}-\sigma}{t}\right)^{2}\right)\sim\left(\frac{Q_{1}+\sigma}{t}\right)^{2}+\left(\frac{Q_{2}-\sigma}{t}\right)^{2}$
and
$\arctan\left(\frac{Q_{2}-\sigma}{t}\right)-\arctan\left(\frac{Q_{1}+\sigma}{t}\right)\sim\frac{Q_{2}-Q_{1}-2\sigma}{t}.$
Then for $\sigma\in[a,b]$ we will have
$\log|t|<|G_{Q_{1},Q_{2}}(\sigma+it)|<(1+O_{Q_{1},Q_{2},\sigma}(1/t^{2}))\log|t|.$
Moreover, $|G(\sigma+it)|$ will be decreasing in $\sigma$ for $\sigma\in[-Q_{1},(Q_{2}-Q_{1})/2]$ and increasing for $\sigma\in[(Q_{2}-Q_{1})/2,Q_{2}]$ provided $t$ sufficiently large (depending on $Q_{1}$ and $Q_{2}$). Hence with a careful choice of $Q_{1}$ and $Q_{2}$ it can satisfy the conditions of Theorem 7.
For iterated logarithms the situation can now be handled as follows. Notice that for appropriate $Q_{1}$ and $Q_{2}$ we have
$\arg(G_{Q_{1},Q_{2}}(s))<O(\tfrac{1}{t})$
and hence
$|\log(G_{Q_{1},Q_{2}}(s))|=\left(1+O\left(\frac{1}{t^{2}(\log t)^{2}}\right)\right)\log|G_{Q_{1},Q_{2}}(s)|$
This has two applications. One, to introduce $\log\log$ terms we can use in Theorem 7, but also, that functions of the form
$H(s)={\rm exp}((\beta-\alpha)s\log(G_{Q_{1},Q_{2}}(s)))G_{Q_{1},Q_{2}}(s)^{\alpha}$
will have
$|H(\sigma+it)|=(1+O(1))(\log|t|)^{\alpha(1-\sigma)+\beta\sigma}$
so that one could potentially directly use it to prove a theorem like Theorem 7, again, potentially at the expense of introducing a constant.
3. Examples
We have the following bounds on $\zeta(s)$ on the line $1+it$ for $|t|>3$
$|\zeta(1+it)|$ $\leq\log|t|$
$|\zeta(1+it)|$ $\leq\tfrac{1}{2}\log|t|+1.93$
$|\zeta(1+it)|$ $\leq\tfrac{1}{5}\log|t|+44.02$
$|\zeta(1+it)|$ $\leq 1.731\frac{\log|t|}{\log\log|t|}$
$|\zeta(1+it)|$ $\leq 58.096(\log|t|)^{2/3}$
which come from [patel_explicit_2022] (first three), [hiary2023explicit] (forth), and [BELLOTTI2024128249] or [leong2024explicitestimateslogarithmicderivative] (for the final). Additionally it is well known that for any $\eta>1$ we have
$|\zeta(\eta+it)|\leq\zeta(\eta).$
Since for $\sigma>1$ we have
$\left|\frac{\sigma+it-1}{\sigma+it}\right|<1$
all of the above bounds also hold with $\zeta(s)$ replaced by $\frac{s-1}{s}\zeta(s)$.
One may verify numerically, the maximum modulus principal simplifying the check, that for $\sigma\in[1,2]$ and $t\in[0,3]$ that $\left|\frac{s-1}{s}\zeta(s)\right|\leq 1$ with the maximum being taken on at $s=1$. Also, for $\sigma\in[1,2]$ and $t\in[0,30]$ we have $\left|\frac{s-1}{s}\zeta(s)\right|\leq 2.1$ the maximum being near $1+27.7i$.
Now, consider the functions
$G_{1}(s)$ $=\frac{1}{2}\left(\log(e+s)+\log(e+2-s)\right)$
$G_{2}(s)$ $=\frac{1}{4}\left(\log(e+s)+\log(e+2-s)\right)+1.93$
$G_{3}(s)$ $=\frac{1}{10}\left(\log(e+s)+\log(e+2-s)\right)+44.02$
$G_{4}(s)$ $=1.731\frac{\log(e+s)+\log(e+2-s)}{2\log\left(\frac{1}{2}\left(\log(e+s)+\log(e+2-s)\right)\right)}$
$G_{5}(s)$ $=58.096\left(\frac{1}{2}\left(\log(e+s)+\log(e+2-s)\right)\right)^{2/3}.$
We verify numerically that for $\sigma\in[1,2]$ and $t\in[0,3]$ the absolute values of each of these is at least $1$, and for $G_{4}(s)$ it is bounded below by $3$ for $\sigma\in[1,2]$ and $t\in[0,30]$ Moreover, for $t\geq 3$ each of $G_{1}$, $G_{2}$, $G_{3}$, and $G_{5}$ are increasing in $\sigma$ and for $t\geq 30$ we have $G_{4}$ is increasing in $\sigma$. For $\sigma\in[1,2]$ and $|t|\geq 3$ we have
$\log|t|$ $\leq\left|\frac{1}{2}\left(\log(e+s)+\log(e+2-s)\right)\right|$
$\frac{\log|t|}{\log\log|t|}$ $\leq\left|\frac{\log(e+s)+\log(e+2-s)}{2\log\left(\frac{1}{2}\left(\log(e+s)+\log(e+2-s)\right)\right)}\right|$
We may thus apply Theorem 7 to the function $f(s)=\frac{s-1}{s}\zeta(s)$ between $\sigma=1$ and $\sigma=\eta$ for some $\eta\in[1,2]$ (for each $G_{i}$ separately) where the exponents on $G_{i}$ are $\alpha=1$ and $\beta=0$ and as in Corollary 3 we introduce $A=1$ and $B=\zeta(\eta)$.
Notice that for the cases of $G_{1}$, $G_{2}$, $G_{3}$ and $G_{5}$ the conditions for $|t|\leq 3$ is trivially verified as $\zeta(\eta)\geq\zeta(2)>1$. Similarly with $G_{4}$ with $|t|\leq 30$ and $\eta<1.5$ as $\zeta(1.5)>2.1$. To extend to $\eta\in[1.5,2]$ one can use that for $\sigma\in[1.5,2]$ and $t\in[0,30]$ we have $\left|\frac{\sigma+it-1}{\sigma+it}\zeta(s)\right|<\zeta(2)$. We thus conclude that for any $\eta\in[1,2]$ that for $\sigma={\rm Re}(s)\in[0,\eta]$ we have
$\left|{\frac{s-1}{s}\zeta(s)}\right|\leq\zeta(\eta)^{\frac{\sigma-1}{\eta-1}}|G_{i}(s)|^{\frac{\eta-\sigma}{\eta-1}}.$
Now notice
$\frac{\left|\sigma+it\right|}{\left|\sigma-1+it\right|}|G_{1}(\sigma+it)|/\log|t|$
$\frac{\left|\sigma+it\right|}{\left|\sigma-1+it\right|}|G_{2}(\sigma+it)|/(\tfrac{1}{2}\log|t|+1.93)$
$\frac{\left|\sigma+it\right|}{\left|\sigma-1+it\right|}|G_{3}(\sigma+it)|/(\tfrac{1}{5}\log|t|+44.02)$
$\frac{\left|\sigma+it\right|}{\left|\sigma-1+it\right|}|G_{4}(\sigma+it)|/(1.731\frac{\log|t|}{\log\log|t|})$
$\frac{\left|\sigma+it\right|}{\left|\sigma-1+it\right|}|G_{5}(\sigma+it)|/(58.096(\log|t|)^{2/3})$
are all decreasing in $t$ with a limit of $1$. This leads to constants $C_{i}(t_{0})$ where
$\left|{\zeta(\sigma+it)}\right|$ $\leq C_{1}(t_{0})\zeta(\eta)^{\frac{\sigma-1}{\eta-1}}(\log|t|)^{\frac{\eta-\sigma}{\eta-1}}$ (1)
$\left|{\zeta(\sigma+it)}\right|$ $\leq C_{2}(t_{0})\zeta(\eta)^{\frac{\sigma-1}{\eta-1}}(\tfrac{1}{2}\log|t|+1.93)^{\frac{\eta-\sigma}{\eta-1}}$ (2)
$\left|{\zeta(\sigma+it)}\right|$ $\leq C_{3}(t_{0})\zeta(\eta)^{\frac{\sigma-1}{\eta-1}}(\tfrac{1}{5}\log|t|+44.02)^{\frac{\eta-\sigma}{\eta-1}}$ (3)
$\left|{\zeta(\sigma+it)}\right|$ $\leq C_{4}(t_{0})\zeta(\eta)^{\frac{\sigma-1}{\eta-1}}(1.731\frac{\log|t|}{\log\log|t|})^{\frac{\eta-\sigma}{\eta-1}}$ (4)
$\left|{\zeta(\sigma+it)}\right|$ $\leq C_{5}(t_{0})\zeta(\eta)^{\frac{\sigma-1}{\eta-1}}(58.096(\log|t|)^{2/3})^{\frac{\eta-\sigma}{\eta-1}}$ (5)
for all $t>t_{0}$. One verifies that in all cases we have
$\begin{array}[]{ccc}C_{i}(10)<1+4.6\cdot 10^{-2}&C_{i}(10^{2})<1+3.2\cdot 10^{-4}&C_{i}(10^{3})<1+2.6\cdot 10^{-6}\\ C_{i}(10^{4})<1+2.4\cdot 10^{-8}&C_{i}(10^{5})<1+2.2\cdot 10^{-10}&C_{i}(10^{6})<1+2.1\cdot 10^{-12}.\end{array}$
[…]